1. Differentiate each of the following functions: (a) 22--6 (b) (1+) (c) In(x3 + x) (d) (2 – 2)e-2 (f) (In x)(:– 1)ez
Answer #1

1) y = x​​​​​​​​​2 - x-6

differentiate both sides with respect to X

dy/dx =( d(x2)/dx) - (d (x-6)/dx)

dy/dx. = 2x -(-6) (x​​​​​​-6-1​​ )

dy/dx = 2x +6x-7

2) y = (1+x​​​​​​2​​​)4

dy/dx = (d(1+x​​​​​​2 )4 /dx) (d (1+x​​​​​​2 )/dx)

dy/dx = 4(1+x​​​​​​2 )4-1 (2x)

dy/dx = 8x (1+x​​​​​​2 )3

3) y = ln ( x​​​​​​3​​​+x)

dy/dx = (d ( ln (x3 + x) / dx) (d (x3 + x)/dx)

dy/dx = (1/(x​​​​​​3 +x) ) (3x3-1 + 1)

dy/ dx = (3x2+1)/(x​​​​​​3+x)